Please show the steps to complete the problem.

A 1500 kg car is approaching a hill shown in the f
Answer #1
A)Initial Total energy = (0.5*1500*(10^2))+(1500*9.81*5) = 148575 J
Final potential energy = (1500*9.81*10) = 147150 J
Initial Total energy > Final potential energy
So it will have some final K.E. So, yes it can make to the top of the hill by coasting.
B)Initial Total energy = Final total energy
=>(0.5*1500*(10^2))+(1500*9.81*5) = (0.5*1500*(V^2))
=>V = 14.07480018 m/s
Answer #2
A) mv^2/2 =mgh,,,,,,,,,,,,,,,,h=v^2/2g=5.1 m ,,,yes car will make to top of hill,,,,,,,,,,,,,B) mg(10)=mv^2/2=14m/s